OpenStudy (anonymous):

A man bought 3 apples,5 mangoes and 8 bananas for $87.Had he bought 5 apples,7 mangoes and 6 bananas it would cost him $121.By how much would an apple and mango together cost more than a banana?

OpenStudy (anonymous):

Put it in terms of equation and tell me what equations can you form from this??

OpenStudy (anonymous):

3a+5b+8c=87 5a+7b+6c=121

OpenStudy (anonymous):

How to solve it..??

OpenStudy (anonymous):

how to solve this?

OpenStudy (anonymous):

don't know. but you ask and you shall receive! (you asked for equations)

OpenStudy (anonymous):

its easy to solve, first you need to put what you want in equation form as well, eli don't help in this case!

OpenStudy (anonymous):

OMG MY NAME IS ELIZA!!!!!!!!!!! and ok :)

OpenStudy (anonymous):

As per the question you have equation as 5a+7b+6c=121 -----(1) 3a+5b+8c=87 -------(2) (1) -(2) you will get the required answer

OpenStudy (anonymous):

i.e 2a+2m-2b = 34 a+m-b = 17

OpenStudy (anonymous):

correct!.. thats what you need!